# Survey of Graph Neural Network for Internet of Things and NextG Networks

### v1

@@ V GNN for Networking
With the evolving number of wireless devices, the research community is working towards finding ways to improve the network performance to meet the quality of service requirements of the users. To this end, many ML-based solutions have been proposed in an effort to achieve the goal of better network performance. In this section, we will focus on applications of GNN for various networking applications such as Network and Signal Characteristics Prediction (Section V-A), Routing Optimization (Section V-B), Congestion Control (Section V-C), Mobile Edge Computing (Section V-D), Digital Twin Networks (Section V-E) and Unmanned Aerial Networks (Section V-F).
@@ V-A Network and Signal Characteristics Prediction
Prediction applications include but not limited to Wireless LAN (WLAN) performance prediction [196, 197], Received Signal Strength Indicator (RSSI) prediction [198], network traffic prediction [199] and channel prediction [200]. In [196], the authors present a data-driven approach using GCN to predict the performance of next-generation WiFi networks. The designed model can predict the throughput of densely deployed WiFi nodes using Channel Bonding (CB). CB is a technique where two or more channels are combined thereby increasing the throughput between multiple wireless devices. In [197], the authors present HTNet, a specialized Heterogeneous temporal GNN that can extract features from dynamic WLAN deployment to predict its throughput. Through analysis of the unique graph structure of WLAN deployment graphs, HTNets can achieve the lowest prediction error in real-world experiments.
@@
In [198] the authors use a combination of Random Dot Product Graphs (RDPG) and GNNs to address the problem of predicting wireless received signal strength (RSS) of wireless networks. RDPG is a spectral-based embedding method that provides a flexible framework for generating random graphs with various structural properties. Finally, in [199] the authors use a time-evolving graph to formulate the time-evolving nature of the user movements and propose a Graph-based Temporal Convolutional Network (GTCN) to predict the future traffic of each node in the wireless network. The GTCN can effectively learn intra- and inter-time spatial correlation between each node through the node aggregation method as well as can efficiently model the temporal dynamics of mobile traffic trends from different nodes through the temporal convolutional layer. In [200], the authors present an environment-information based channel prediction method to effectively predict the large-scale and small-scale fading parameters using GNN. First, scatterer-centered communication environment graphs is used to classify the scattering sources affecting the channel characteristics into effective, primary, and non-primary scatterers. The scatterers that produce multi-path and influence the channel characteristics are referred to as effective scatterers while the scatters that produce single propagation points are referred to as primary scatterers and all other scatterers are non-primary scatterers. After successful classification, the GNN-based channel prediction module accomplishes the task of predicting the large-scale and small-scale fading parameters using fully-connected propagation environment graphs and single-connected propagation environment graphs. Furthermore, a GNN-based multi-target regression model is designed for continuous channel parameters prediction.
@@ V-B Routing Optimization
The key role of the network layer; the third layer of the Open Systems Interconnection (OSI) model is routing packets through the network. There are various routing protocols designed to facilitate the exchange of routing information among routers and enable them to make informed decisions about packet forwarding. The objective of routing protocols is to determine the best path or route for data packets from a source to a destination in a network based on predefined performance metrics (e.g. minimizing latency or minimizing energy consumption). Thus it is important to have an optimized routing algorithm that can improve efficiency by selecting the best routes, thereby reducing network congestion and improving overall network performance, resulting in better Resource Utilization.
Traditionally, techniques such as reinforcement learning, and recurrent neural networks were commonly used for routing optimization. Recently researchers have studied the applicability of GNN for routing optimization due to its ability to model and learn from graph-structured data. The key advantage of GNN compared to other methods is that by representing the network topology as a graph, where nodes represent routers or switches and edges represent links, GNNs can learn to predict the best paths for forwarding packets based on factors such as link capacities, congestion levels, latency and much more. The application of GNN for routing optimization has gained significant traction in the research community [122, 135, 201, 202, 203, 204, 205].
@@
For example, in [122], the authors use a GCN-based DRL algorithm to jointly optimize traffic scheduling and routing. The key advantage of this algorithm is that there is no prior knowledge of the system, and all the information is obtained in run-time by interacting with the environment. In [135], they propose a reliable routing optimization algorithm using GraphSAGE. The algorithm consists of two components. First, a multi-path routing algorithm using breadth-first search method that can continuously predict the shortest path for different network topologies. Second is GraphSAGE-based routing performance optimization algorithm that creates network labels for training GraphSAGE and are used to learn graph features to predict the network performance for each path and select the best path for optimal network performance. This proposed method can minimize network congestion, improve network throughput, and reduce end-to-end delay and packet loss rate compared to traditional shortest-path routing algorithms.
RouteNet [201] is a novel network model-based GNN that can understand the complex relationships between topology, routing and input traffic and accurately predict the estimate of per-source/destination per-packet delay and packet loss. RouteNet can be generalized to arbitrary topologies, routing schemes, and traffic intensity thanks to GNN’s ability to learn and model graph-structured information. In [202] the authors explore the advantage of integrating GNN with Deep Reinforcement learning (DRL) to overcome the challenges of using DRL which are not well suited to learning from information structured as graphs and the limited generalizability of DRL. The proposed GNN-based DRL algorithm can learn and generalize to any given network topology. The proposed algorithm not only works for scenarios seen during the training process but also for network topologies that were not part of the training set. Similarly, in [203], the authors propose GRL-NET, a combination of GNN and DRL-based intelligent routing algorithm. Compared to using conventional methods to construct DRL agent, in this work, GNN is used due to its ability to learn from complex relationships among network topologies.
@@ V-C Congestion control
Congestion control refers to the techniques and mechanisms used in computer networks to regulate the flow of data packets, preventing network congestion, and ensuring efficient data transmission. When network resources such as bandwidth, buffers, or processing capacity are overwhelmed with data, congestion occurs, leading to packet loss, increased latency, and reduced throughput. Congestion control aims to manage this situation to maintain network stability and performance. While traditional ML algorithms are extensively used for congestion control, now the research community is investigating the use of GNN to efficiently control the congestion in small and large-scale networks. Some of the research work includes CongestionNet [206], CONAIR [207] and RouteNet-Fermi [208].
@@ V-D Mobile Edge Computing
Mobile Edge Computing (MEC) is a decentralized computing paradigm that brings computation and data storage closer to the location where it is needed, rather than relying solely on centralized cloud data centers. In edge computing, data is processed and analyzed locally on edge devices or servers, often at or near the data source, such as IoT devices, sensors, and other endpoints. Edge computing and GNNs can be combined to create intelligent, real-time decision-making systems at the edge of the network. This combination is particularly powerful for applications that involve analyzing data from a network of interconnected devices or sensors, where GNNs can capture complex relationships and patterns in the data. Figure 9 shows the workflow of GNN-enabled MEC. The workflow can be divided into six steps. First, the user equipment (such as a cellular device or cellular-enabled UAV) sends the computation/task request and relevant information (CTRI) to the base station. Second, the base stations aggregate all the task information and QoS requirements and forwards the combined combined CTRI to the Edge Server Controller (ESC). Third, The ESC collects the information, generates the graph representation and the GNN decision agent (usually coupled with an RL) decides the best edge server capable of executing the task that satisfies the user QoS and sends a computation/task allocation (CTA) message to the edge server. Fourth, The BS then relays this CTA message to the user. Fifth, the user sends the task data to the edge server through the BS. Sixth, the edge server executes the task and sends the computation/task output (CTO) back to the user. Each of these steps is clearly labeled with numbered blue circles for cellular-enabled UAVs, numbered red circles for cellular users, and numbered purple circles for the combined messages.
@@
Task offloading in MEC refers to the process of transferring computational tasks from mobile devices (such as smartphones, tablets, or IoT devices) to nearby edge servers or nodes located at the edge of the network infrastructure. In traditional cloud computing models, all computation is typically performed in centralized data centers, which can result in high latency and network congestion, especially for latency-sensitive applications running on mobile devices. Task offloading in MEC aims to alleviate these issues by bringing computation closer to the end-users, reducing latency, and improving the overall quality of service. GNN and Task offloading have been extensively studied and are applied to various applications [212, 213, 214, 215, 216]. For example, in [212] the authors propose a task offloading mechanism based on graph neural network called graph reinforcement learning-based offloading (GRLO) framework is paired with an actor-critic network to overcome the challenges of Heuristic algorithms and deep reinforcement learning-based approaches that heavily depend on accurate mathematical models for the MEC system, and deep reinforcement learning techniques does not make fair use of the relationship between devices in the network. Similarly, in [213], the authors propose a meta-reinforcement learning task offloading algorithm called GASTO to overcome the challenges associated with the generalization and robustness of offloading algorithms in dynamic and uncertainties in the real-world environment. In [214], the authors propose a task offloading mechanism based on GCN which generates the graph and analyzes the integrated relationship from features such as task characteristics, network conditions, and available resources at the edge. GCN enables improved decision-making, enhanced resource utilization, and optimized performance in the edge network. In [215] the authors present an end-to-end fine-grained computing offloading model with the goal of achieving better load balancing. To this end, a deep graph matching method based on GNN is used to select the best node for offloading and can be applied for dynamic and large-scale networks and achieve faster execution speed. In comparison to existing methods for task offloading and load balancing, the proposed method can reduce the network load imbalance efficiently and due to its shorter execution time, the proposed scheme is suitable for real-time scenarios with multiple subtasks and edge nodes. Similarly, in [216], the authors propose a GCN-RL based algorithm for Container Scheduling for NextG IoT user task offloading and resource scheduling architecture with the objective of minimizing average task completion time for fine-grained and dynamic container scheduling problems.
@@ V-E Digital Twin Networks
One of the most promising enabling techniques for future IoT and next-generation networks is Digital Twins (DT) [217]. Digital Twins are a virtual replica of the real world which enables researchers to create an accurate model of the real-world environment to test and improve their ML algorithms before integrating with real-world hardware. Figure 10 shows an example workflow diagram of GNN-enabled DT network.
@@
For example, in [220], the authors present a scalable DT of 5G network to model the interactions between multiple network slices into a graph and use a GNN-based virtual representation model that captures the hidden interdependencies among nodes based on the aggregation of information from multiple neighbors. Here, the DT is used to explore the complex traffic generated by the network slices and accurately predict the end-to-end slice latency. The DT designed based on GNN is shown to have higher accuracy in estimating the end-to-end latency for different topologies, traffic distributions, and QoS constraints that were not part of the training process. Similarly, in [221], the authors propose a DT-enabled deep distributed Q-network (DDQN) framework that constructs a digital replica of physical slicing-enabled beyond 5G networks that can simulate complex environment and predict its dynamic characteristics. Here, the DT is represented using graph and GNN to learn the complicated relationships of the network slice and the network states are forwarded to the DDQN agent to learn the optimal network slicing policy. It is shown that the proposed DT model achieves optimal and autonomous network management of complex slicing-enabled beyond 5G networks.

### v2

@@ V GNN for Networking
With the evolving number of wireless devices, the research community is working towards finding ways to improve the network performance to meet the quality of service requirements of the users. Many ML-based solutions have been proposed in an effort to achieve the goal of better network performance. GNNs have gained significant attention due to their ability to model complex relationships in network structures and has been successfully applied to various networking challenges, offering promising improvements in performance and decision-making capabilities. In this section we aim to provide a comprehensive overview of how GNNs are shaping the future of networking to create more intelligent, efficient, and adaptive networking infrastructures. To this end, we will focus on applications of GNN for various networking challenges such as Network and Signal Characteristics Prediction (Section V-A), Routing Optimization (Section V-B), Congestion Control (Section V-C). We will discuss the application of GNN for Mobile Edge Computing (Section V-D). Within the networking domain, there are two rapidly growing fields of research namely Digital Twin Networks and Unmanned Aerial Networks. In this section, we will discuss the application of GNN for Digital Twin Networks in Section V-E and Unmanned Aerial Networks in Section V-F. Table VII provides a high-level comparison of surveyed paper using GNNs for wireless networking applications. We also provide the dataset or the simulator used for performance evaluation.
@@ V-A Network and Signal Characteristics Prediction
Prediction applications include but not limited to Wireless LAN (WLAN) performance prediction [233, 236], Received Signal Strength Indicator (RSSI) prediction [237], network traffic prediction [239], channel prediction [240] and link prediction [243]. In [233], the authors present a data-driven approach using GCN to predict the performance of next-generation WiFi networks. The designed model can predict the throughput of densely deployed WiFi nodes using Channel Bonding (CB). CB is a technique where two or more channels are combined thereby increasing the throughput between multiple wireless devices. In [236], the authors present HTNet, a specialized Heterogeneous temporal GNN that can extract features from dynamic WLAN deployment to predict its throughput. Through analysis of the unique graph structure of WLAN deployment graphs, HTNets can achieve the lowest prediction error in real-world experiments.
@@
In [237] the authors use a combination of Random Dot Product Graphs (RDPG) and GNNs to address the problem of predicting wireless received signal strength (RSS) of wireless networks. RDPG is a spectral-based embedding method that provides a flexible framework for generating random graphs with various structural properties. Finally, in [239] the authors use a time-evolving graph to formulate the time-evolving nature of the user movements and propose a Graph-based Temporal Convolutional Network (GTCN) to predict the future traffic of each node in the wireless network. The GTCN can effectively learn intra- and inter-time spatial correlation between each node through the node aggregation method as well as can efficiently model the temporal dynamics of mobile traffic trends from different nodes through the temporal convolutional layer. In [240], the authors present an environment-information based channel prediction method to effectively predict the large-scale and small-scale fading parameters using GNN. First, scatterer-centered communication environment graphs are used to classify the scattering sources affecting the channel characteristics into effective, primary, and non-primary scatterers. The scatterers that produce multi-path and influence the channel characteristics are referred to as effective scatterers while the scatters that produce single propagation points are referred to as primary scatterers and all other scatterers are non-primary scatterers. After successful classification, the GNN-based channel prediction module accomplishes the task of predicting the large-scale and small-scale fading parameters using fully-connected propagation environment graphs and single-connected propagation environment graphs. Furthermore, a GNN-based multi-target regression model is designed for continuous channel parameters prediction. In [243], the authors propose a link prediction approach using GCapsNet called Graph Conversion Capsule Link (GCCL) which is capable of operating on graphs with and without node features. This is achieved by first using GNN to generate the node embeddings and then transforming the node embedding to a node pair feature map. Then CapsNet is applied to learn the feature representation of the feature map. Through this transformation, the link prediction problem is similar to a binary classification problem where the two decisions are either the two nodes are likely to have an edge in the near future or not.
@@ V-B Routing Optimization
In the previous section we discussed the applications of GNN for network and signal characteristics prediction. The next step is to use the predicted characteristics and decide on the optimal routing path to further improve the network performance. To this end, in this section, we will briefly introduce the process of routing and then delve into the applications of GNN for routing optimization. There are various routing protocols designed to facilitate the exchange of routing information among routers and enable them to make informed decisions about packet forwarding. The objective of routing protocols is to determine the best path or route for data packets from a source to a destination in a network based on predefined performance metrics (e.g. minimizing latency or minimizing energy consumption). Thus it is important to have an optimized routing algorithm that can improve efficiency by selecting the best routes, thereby reducing network congestion and improving overall network performance, resulting in better Resource Utilization.
Traditionally, techniques such as reinforcement learning, and recurrent neural networks were commonly used for routing optimization. Recently researchers have studied the applicability of GNN for routing optimization due to its ability to model and learn from graph-structured data. The key advantage of GNN compared to other methods is that by representing the network topology as a graph, where nodes represent routers or switches and edges represent links, GNNs can learn to predict the best paths for forwarding packets based on factors such as link capacities, congestion levels, latency and much more. The application of GNN for routing optimization has gained significant traction in the research community [134, 147, 246, 248, 250, 251, 252, 283].
@@
For example, in [134], the authors use a GCN-based DRL algorithm to jointly optimize traffic scheduling and routing. The key advantage of this algorithm is that there is no prior knowledge of the system, and all the information is obtained in run-time by interacting with the environment. In [147], they propose a reliable routing optimization algorithm using GraphSAGE. The algorithm consists of two components. First, a multi-path routing algorithm using breadth-first search method that can continuously predict the shortest path for different network topologies. Second is GraphSAGE-based routing performance optimization algorithm that creates network labels for training GraphSAGE and is used to learn graph features to predict the network performance for each path and select the best path for optimal network performance. This proposed method can minimize network congestion, improve network throughput, and reduce end-to-end delay and packet loss rate compared to traditional shortest-path routing algorithms.
RouteNet [246] is a novel network model-based GNN that can understand the complex relationships between topology, routing and input traffic and accurately predict the estimate of per-source/destination per-packet delay and packet loss. RouteNet can be generalized to arbitrary topologies, routing schemes, and traffic intensity thanks to GNN’s ability to learn and model graph-structured information. In [248] the authors explore the advantage of integrating GNN with Deep Reinforcement learning (DRL) to overcome the challenges of using DRL which is not well suited to learning from information structured as graphs and the limited generalizability of DRL. The proposed GNN-based DRL algorithm can learn and generalize to any given network topology. The proposed algorithm not only works for scenarios seen during the training process but also for network topologies that were not part of the training set. Similarly, in [250], the authors propose GRL-NET, a combination of GNN and DRL-based intelligent routing algorithm. Compared to using conventional methods to construct DRL agent, in this work, GNN is used due to its ability to learn from complex relationships among network topologies.
@@ V-C Congestion Control
While routing optimization tries to improve the network performance, as more and more devices join the network and all devices request data with specific QoS requirements, it is important to serve all the users simultaneously which could cause congestion in the network. Even though routing optimization finds the optimal path, it does not necessarily control the flow of data within the route. Therefore it is important to control the data flow which is called congestion control. In this section, we first introduce congestion control and then look into the application of GNN to address this challenge. Congestion control refers to the techniques and mechanisms used in computer networks to regulate the flow of data packets, preventing network congestion, and ensuring efficient data transmission. When network resources such as bandwidth, buffers, or processing capacity are overwhelmed with data, congestion occurs, leading to packet loss, increased latency, and reduced throughput. Congestion control aims to manage this situation to maintain network stability and performance. While traditional ML algorithms are extensively used for congestion control, now the research community is investigating the use of GNN to efficiently control the congestion in small and large-scale networks. Some of the research work includes CongestionNet [253], CONAIR [254] and RouteNet-Fermi [256].
@@ V-D Mobile Edge Computing
In the previous section we mentioned that devices are constantly gathering information and need to process them to extract the relevant information and take necessary actions to achieve its goal. While large capable types of machinery can process huge amounts of data, machinery that are battery-constrained, low-power with limited computation and storage devices cannot effectively process and store the data. This raises the need for an effective computing and storage mechanism that can be leveraged by such devices. One of the common solutions to this challenge is called Mobile Edge Computing. In this section, we will briefly introduce mobile edge computing before discussing how GNN is applied in various applications. Mobile Edge Computing (MEC) is a decentralized computing paradigm that brings computation and data storage closer to the location where it is needed, rather than relying solely on centralized cloud data centers. In edge computing, data is processed and analyzed locally on edge devices or servers, often at or near the data source, such as IoT devices, sensors, and other endpoints. Graphs are commonly used to represent the network topology and find its applications in Multi-Edge Cooperative Computing [284] and Edge Cloud system for prioritized services [285], to name a few where edge computing and GNNs can be combined to create intelligent, real-time decision-making systems at the edge of the network. This combination is particularly powerful for applications that involve analyzing data from a network of interconnected devices or sensors, where GNNs can capture complex relationships and patterns in the data. Figure 12 shows the workflow of GNN-enabled MEC. The workflow can be divided into six steps. First, the user equipment (such as a cellular device or cellular-enabled UAV) sends the computation/task request and relevant information (CTRI) to the base station. Second, the base stations aggregate all the task information and QoS requirements and forwards the combined CTRI to the Edge Server Controller (ESC). Third, The ESC collects the information, generates the graph representation and the GNN decision agent (usually coupled with an RL) decides the best edge server capable of executing the task that satisfies the user QoS and sends a computation/task allocation (CTA) message to the edge server. Fourth, The BS then relays this CTA message to the user. Fifth, the user sends the task data to the edge server through the BS. Sixth, the edge server executes the task and sends the computation/task output (CTO) back to the user. Each of these steps is clearly labeled with numbered blue circles for cellular-enabled UAVs, numbered red circles for cellular users, and numbered purple circles for the combined messages.
@@
Task offloading in MEC refers to the process of transferring computational tasks from mobile devices (such as smartphones, tablets, or IoT devices) to nearby edge servers or nodes located at the edge of the network infrastructure. In traditional cloud computing models, all computation is typically performed in centralized data centers, which can result in high latency and network congestion, especially for latency-sensitive applications running on mobile devices. Task offloading in MEC aims to alleviate these issues by bringing computation closer to the end-users, reducing latency, and improving the overall quality of service. GNN and Task offloading have been extensively studied and are applied to various applications [177, 264, 124, 265, 266]. For example, in [177] the authors propose a task offloading mechanism based on graph neural network called graph reinforcement learning-based offloading (GRLO) framework is paired with an actor-critic network to overcome the challenges of Heuristic algorithms and deep reinforcement learning-based approaches that heavily depend on accurate mathematical models for the MEC system, and deep reinforcement learning techniques does not make fair use of the relationship between devices in the network. Similarly, in [264], the authors propose a meta-reinforcement learning task offloading algorithm called GASTO to overcome the challenges associated with the generalization and robustness of offloading algorithms in dynamic and uncertainties in the real-world environment. In [124], the authors propose a task offloading mechanism based on GCN which generates the graph and analyzes the integrated relationship from features such as task characteristics, network conditions, and available resources at the edge. GCN enables improved decision-making, enhanced resource utilization, and optimized performance in the edge network. In [265] the authors present an end-to-end fine-grained computing offloading model with the goal of achieving better load balancing. To this end, a deep graph matching method based on GNN is used to select the best node for offloading and can be applied for dynamic and large-scale networks and achieve faster execution speed. In comparison to existing methods for task offloading and load balancing, the proposed method can reduce the network load imbalance efficiently and due to its shorter execution time, the proposed scheme is suitable for real-time scenarios with multiple sub-tasks and edge nodes. Similarly, in [266], the authors propose a GCN-RL based algorithm for Container Scheduling for NextG IoT user task offloading and resource scheduling architecture with the objective of minimizing average task completion time for fine-grained and dynamic container scheduling problems.
@@ V-E Digital Twin Networks
In prior sections we discussed the common challenges and applications that used GNN that are not necessarily exclusive to NextG networks, but they are still important for any wireless networks. To shed some light on some NextG applications and the applications of GNN, in this section, we discuss one of the exciting and extensively researched topics by the community called Digital Twins. Digital Twins (DT) is considered one of the most promising enabling techniques for future IoT and next-generation networks [286]. Digital Twins are a virtual replica of the real world which enables researchers to create an accurate model of the real-world environment to test and improve their ML algorithms before integrating with real-world hardware. Figure 13 shows an example workflow diagram of the GNN-enabled DT network.
@@
For example, in [267], the authors present a scalable DT of 5G network to model the interactions between multiple network slices into a graph and use a GNN-based virtual representation model that captures the hidden inter-dependencies among nodes based on the aggregation of information from multiple neighbors. Here, the DT is used to explore the complex traffic generated by the network slices and accurately predict the end-to-end slice latency. The DT designed based on GNN is shown to have higher accuracy in estimating the end-to-end latency for different topologies, traffic distributions, and QoS constraints that were not part of the training process. Similarly, in [269], the authors propose a DT-enabled deep distributed Q-network (DDQN) framework that constructs a digital replica of physical slicing-enabled beyond 5G networks that can simulate complex environment and predict its dynamic characteristics. Here, the DT is represented using graph and GNN to learn the complicated relationships of the network slice and the network states are forwarded to the DDQN agent to learn the optimal network slicing policy. It is shown that the proposed DT model achieves optimal and autonomous network management of complex slicing-enabled beyond 5G networks.
In [270], the authors propose TwinNet, a GNN-based DT model that can take into account the complex relationship between different service level agreements metrics such as queuing policy, network topology, routing configuration, and input traffic. It is shown that GNN-enabled TwinNet can generalize to its input parameters, topologies, routing and queuing configuration, and can accurately predict the end-to-end path delays in 106 unseen real-world topologies from Internet Topology Zoo [289]. Flow-aware Digital Twin (FlowDT) [272], is a GNN-based framework to model computer networks, where flow refers to the set of packets that share common characteristics such as the same protocols (UDP/TCP), IP addresses or port numbers of source and destination node. FlowDT supports the creation and destruction of flows and also provides accurate estimates of per-flow metrics such as delay and jitter.
@@
V-F Unmanned Aerial Networks
While traditional base stations offer good network coverage over longer distances, there arise situations where the existing infrastructure may be damaged due to unforeseen circumstances. This has motivated researchers to find a solution that can serve as a temporary solution while the existing infrastructure is brought back to working order. This has given rise to Unmanned Aerial Networks. In this section, we will first discuss unmanned aerial networks and its applications and then discuss how GNNs are used to address some of the challenges. Unmanned Aerial Networks (UANs) refer to networks formed by interconnected Unmanned Aerial Vehicles (UAVs), also known as drones, operating cooperatively to achieve various objectives. Due to the portable nature of the UAVs, they can be dynamically deployed when required and play a crucial role in establishing temporary connectivity in the event of disaster or infrastructure failure and enable high data rate network connectivity. Figures 14 and 15 show the Typical components of UANs and applications of GNN for UAN tasks, respectively. ML techniques such as RNNs and RL are commonly used for UAN problems such as spectrum access and flight control for microwave, millimeter wave, and Terahertz frequency bands [290, 291, 292, 293, 294, 295], UAV swarm collaborative beamforming [296, 297, 298, 299]. In addition to traditional ML techniques GNNs are also commonly used for UANs problems such as Sensing [275] and networking applications such as trajectory planning [277, 279], resilient UAV swarm communication [280], Flying Relay Location and Routing Optimization [282].
In [275], the authors propose a GCN and GAT-based LSTM model to accurately predict the real-time and future Air Quality Index (AQI) using aerial-ground sensing framework. This model uses a lightweight Dense-MobileNet model [300] to achieve energy-efficient end-to-end learning from features extracted from images captured by the UAV. This framework allows for federation across various institutions to collaboratively learn a well-trained global model to accurately and securely monitor the AQI and also expand the scope of UAV swarm monitoring. In [277], the authors present a GCN-based framework to solve the cooperative trajectory design problem formulated based on modeling the relationship between the ground terminal and UAV-base stations (UAV-BS). This framework enables UAV-BS to efficiently manage the time-varying local observation such as current location, average throughput, and scheduling and facilitate cooperation between UAV-BSs through information exchange. Due to the lack of global information and a globally optimal solution, a Q-learning-based multi-agent RL algorithm is used by the UAV-BS to learn a distributed trajectory policy which is modeled as a Markov game. Here, the message passing by the GCN framework ensures cooperative behavior among UAV-BS and significantly enhances the network’s performance. In [279], the authors propose a multi-dimensional trajectory prediction model for UAV swarms by integrating GNN with a dynamic graph neural network (DynGN). The DynGN uses an encode-decoder structure with a GCN to process the dynamic adjacency matrix, extract spatial features, and gated recurrent unit to model the temporal dynamics and capture the evolution of UAV interactions over time. This model processes both the dynamically changing network configuration and the trajectory data of the UAV swarm simultaneously.
In [280], the authors study the self-healing problem of UAV swarm network to re-establish network connectivity in the event of unpredictable destruction. These destruction may affect one or more UAVs simultaneously and can either be one-off events or recurring events. To address the one-off events, a GCN-based framework is designed to find the recovery topology of the UAV swarm network in a timely manner. This framework is then extended by incorporating a monitoring mechanism to detect UAV destruction and a self-healing trajectory planning algorithm based on GCN to address the recurring events of destruction. Finally, a meta-learning scheme is designed for the GCN to reduce the execution time complexity and ensure faster convergence of GCN. In [282], the authors propose a GNN-based framework for the optimization of UAV locations and relay paths in UAV-relayed 6G IoT networks. A two-stage GNN model is developed to decouple the problem of optimal relay path selection and UAV position optimization and is solved by a separate GNN model. Here, both stages are trained using reinforcement and unsupervised learning with no training data or prior knowledge of optimal solutions. First, the locations of the UAVs are optimized and second, the best relay path is determined based on the optimized location of the UAVs from the first step. Results show that the proposed model performs better than brute search with lower time complexity, scalable to very large networks with dynamically changing environments.